OpenStudy (anonymous):

Scores on a standardized test are normally distributed with a mean of 350 and a standard deviation of 25. Approximately what percent of students scored between 360 and 380? 10% 16% 23% 38% Question 9.9. In a standard normal distribution, above which z-score do approximately 57% of the data lie? 1.72 0.18 0.18 1.72

OpenStudy (anonymous):

@Kitten_is_back

OpenStudy (anonymous):

@chlobohoe

OpenStudy (kitten_is_back):

which one do u think is your andwer

OpenStudy (anonymous):

okay so do you know the formula for this babe?

OpenStudy (anonymous):

I was thinking number 1 was C and number 2 is D

OpenStudy (anonymous):

im pretty sure youre right about number 1 being C.. do you know how you got that?

OpenStudy (anonymous):

@chlobohoe I'm sorry I really don't know I took a guess for it

OpenStudy (anonymous):

okay so we know two things.. μ = 350 σ = 25 dont let the fancy letters freak you out tho

OpenStudy (anonymous):

the formula should look something like this.. \[z=\frac{ x-\mu }{ \sigma } \]

OpenStudy (anonymous):

okay yes 350 is the mean and 25 is the standard deviation

OpenStudy (anonymous):

okay how do I find the percent to that

OpenStudy (anonymous):

My answer is 14.4,15.2

OpenStudy (anonymous):

standardize x to z = (x - μ) / σ P( 360 < x < 380) = P[( 360 - 350) / 25 < Z < ( 380 - 350) / 25] P( 0.4 < Z < 1.2) = P( z < 1.2) - P( z < 0.4) = 0.8849 - 0.6554 = 0.2295 = 23 %

OpenStudy (anonymous):

and if ur wondering how we changed the original equation to what we did z(360) = 10/25 = .4 z(380) = 30/25 = 1.2 P(350 < x < 380) = normalcdf(.4, 1.2)

OpenStudy (anonymous):

Hmmmm okay I think I got it a little

OpenStudy (anonymous):

Thank you